# TituLLMs: A Family of Bangla LLMs with Comprehensive Benchmarking

### v1

@@ Abstract
In this paper, we present TituLLMs, the first large pretrained Bangla LLMs, available in 1b and 3b parameter sizes. Due to computational constraints during both training and inference, we focused on smaller models. To train TituLLMs, we collected a pretraining dataset of approximately $\sim 37$ billion tokens. We extended the Llama-3.2 tokenizer to incorporate language- and culture-specific knowledge, which also enables faster training and inference. There was a lack of benchmarking datasets to benchmark LLMs for Bangla. To address this gap, we developed five benchmarking datasets. We benchmarked various LLMs, including TituLLMs, and demonstrated that TituLLMs outperforms its initial multilingual versions. However, this is not always the case, highlighting the complexities of language adaptation. Our work lays the groundwork for adapting existing multilingual open models to other low-resource languages. To facilitate broader adoption and further research, we have made the TituLLMs models and benchmarking datasets publicly available.111anonymous.com
@@
Shahriar Kabir Nahin1, Rabindra Nath Nandi1, Sagor Sarker1, Quazi Sarwar Muhtaseem1, Md Kowsher2, Apu Chandraw Shill1, Md Ibrahim1, Mehadi Hasan Menon1, Tareq Al Muntasir1, Firoj Alam3, 1Hishab Singapore Pte. Ltd, Singapore, 2University of Central Florida, USA 3Qatar Computing Research Institute, Qatar
@@ 1 Introduction
The rapid advancements in large language models (LLMs) have reshaped the field of artificial intelligence, showcasing remarkable versatility across numerous tasks Brown et al. (2020); Ouyang et al. (2022); Achiam et al. (2023); Chowdhery et al. (2023). These models demonstrate not only an ability to perform various NLP tasks but also an intriguing potential for self-assessment and continuous improvement Liu et al. (2023b); Fu et al. (2023); Chiang et al. (2023).
@@
  • We proposed a novel translation techniques that helps to develop high quality benchmarking dataset.
@@ 2 Pretraining Data
Pretraining data for Bangla is very limited compared to very high quality data available for English and other high resource languages Penedo et al. (2023); Soldaini et al. (2024). Hence, we needed collect pretraining dataset for training TituLLMs. We have compiled a substantial Bangla raw dataset from a diverse range of sources, encompassing both formal and informal linguistic styles. The dataset is primarily derived from three key sources: web documents, books, and synthetically generated text. The synthetically generated data includes translated data, transliterated data, audio transcripted data, etc. An outline of our data collection and preprocessing pipeline is show in Figure 5. The final high-quality dataset amounts to approximately 268 GB, with 22 GB designated as a validation set, sampled proportionally to maintain the heterogeneity of the data. In total, the corpus contains $\sim 37$ billion tokens, optimized for training and evaluation across various Bangla language processing applications. In Table 3 (in Appendix) and in Figure 3, we report the distribution of tokens for different sources.
@@ 2.3 Synthetic Data
For training the en-bn machine translation (MT) model, we collected open-source translation data from various platforms, including BanglaNMT Hasan et al. (2020a) and Samanantar Ramesh et al. (2022). Furthermore, we generated synthetic bn-en translation pairs using Bangla news sources and Wikipedia as source data, employing Llama-3.1-8B-Instruct  Touvron et al. (2023) for target data generation. We selected Llama for this task due to its superior English-language capabilities.
Using this approach, we created a dataset comprising approximately 60 million translation pairs, which we then used to train the NLLB model. On our in-house test dataset, the BLEU score of this model is 37.6. Once the model has been trained, we have used it to translate a corpus of English news articles  [^5] https://www.kaggle.com/datasets/davidmckinley/all-the-news-dataset into Bangla.
@@
The use of romanized text is very common in everyday communication for Bangla Fahim et al. (2024). To address this, we have developed a Bangla-to-Romanized Bangla dataset by training an NLLB-based (600M-Distilled) model. For model training, we collected transliteration pairs from the Sangraha dataset and generated additional synthetic transliteration pairs using the GPT-4 model Achiam et al. (2023). We then used this dataset to train the NLLB-based transliteration model. The BLEU score for this model is 65.1, as evaluated on an in-house test dataset. We then used this model to create the transliteration dataset by selecting a small subset of collected Bangla Wikipedia articles.
@@
To enhance the model with conversational capabilities, we enriched our dataset by incorporating conversational data. We have crawled topics (e.g., “Rabindranath Tagore’s contributions to Bengali art”) from Wikipedia and Banglapedia on which we generated conversations between two agents. To achieve this we have developed an agentic system where two agents interacted each other on a given topic. In Table 6, we have provided examples of topic, roles and a detail of the prompt. An example conversation is provide in Figure 5. The average number of turns per conversation is 8. In total, we added approximately 1 million conversational data to the dataset.
@@ 2.4 Sangraha Dataset
Additionally, we enriched our dataset by integrating the open-source Sangraha dataset Khan et al. (2024). Sangraha is the largest high-quality, cleaned Indic language corpus. We incorporated the Bangla portion of the Sangraha dataset into our training set.
@@ 3.1 Tokenizer Training
We developed a custom tokenizer for Bangla text using Tiktoken [^6] https://github.com/openai/tiktoken, which employs Byte Pair Encoding (BPE) Sennrich (2015) for tokenization. To train this tokenizer, we sampled 48 GB of data from our pretraining Bangla corpus. Additionally, we modified the original Tiktoken codebase to enhance its efficiency and better accommodate the morphological complexities of Bangla. After training multiple tokenizers on the same subset, we merged each newly trained tokenizer with the existing Llama-3.2 tokenizer. This merging process aimed to preserve the strengths of the original tokenizer while integrating domain-specific vocabulary and improving segmentation for Bangla. To evaluate the performance of each merged tokenizer, we computed the average tokens per word (TPW) on a separate 1 GB sample from the original corpus. Table 4, in Appendix, summarizes the TPW values for both the original Llama-3.2 tokenizer and the newly trained tokenizers. A lower TPW generally indicates more efficient segmentation, which reduces sequence lengths and may enhance downstream model performance.
@@ 3.3 Pretraining
After modifying the models, we pre-trained them on our full dataset using LlamaFactory Zheng et al. (2024). Both models were trained with a context length of 4096, with packing enabled for maximum efficiency. Training for one epoch required 1750 H100 GPU hours.
@@ 4.2 Banchmarking Datasets
We benchmarked TituLLMs alongside other popular LLMs using five newly prepared evaluation datasets. Below, we describe the development process for each dataset. Table 1 presents the distribution and splits of each dataset.
Bangla MMLU: We curated multiple-choice questions from various opensource educational websites and textbooks, inspired by the original MMLU dataset Hendrycks et al. (2020). The dataset includes multiple-choice questions from different Bangladeshi exams, such as job exams, the Bangladesh Civil Service Exam, and undergraduate admission exams. In Figure 7, we report category wise distributions.
CommonsenseQA Bangla (CSQA): We translated the CommonsenseQA dataset Talmor et al. (2018) into Bangla using our custom translation-based approach, Expressive Semantic Translation (EST). This method generates multiple translations for a sentence and iteratively refines them to select the most suitable version. More details on this approach are discussed in Section C.
@@
Bangla MMLU Manual - 72,944 14750 175
@@ 5 Results and Discussion
BoolQ: The BoolQ dataset measures the performance of the model for yes/no question-answering in Bangla. TituLLM-1b achieves 0.53 in the 0-shot setting but drops slightly to 0.51 in the 5-shot setting. In contrast, TituLLM-3B moves from 0.53 (0-shot) to 0.54 (5-shot). However, Llama-3b and Qwen-2.5-1.5b have done much better in this task. As the context length for all BoolQ data was as large as a News document, TituLLM’s performance may drop for long contexts. It suggests further pertaining to the model should be done with long contexts Chowdhery et al. (2022); Kaplan et al. (2020).
@@
Performance of tokenizer: The superior performance of our models in reasoning tasks is mainly an impact of our extended tokenizer. To justify this, we can observe the results of the BongLLaMa models. These models are continual pretrained models with existing open-source Bangla text corpus. If only the dataset could improve the performance then that would be reflected in BongLLaMA models. But we can see they are performing similarly to Llama models. To have an interpretation of our extended tokenizer’s performance we can look into Figure-4. The figure shows Llama tokens and TituLLM tokens for a simple sentence in Bangla with two of the most common words. We can see that Llama tokenizer splits the text into character and byte levels. On the other hand, TituLLM tokenizes the sentence into word or subword levels. As a result, TituLLM can deliver more meaningful tokens than Llama for Bangla text. This is an important advantage of TituLLM that not only enables TituLLM to perform better with smaller datasets but also ensures low latency during inference.
@@ 6 Related Work
The evolution of token adaptation in NLP has progressed from linguistic cues and statistical methods Creutz and Lagus (2006); Luong et al. (2013); Zhang et al. (2023) to phrase-level segmentation Koehn et al. (2007, 2003). The rise of deep learning shifted the focus to subword-level segmentation, enhancing the handling of rare words Sennrich (2015); Kudo (2018); Kudo and Richardson (2018). More recent efforts emphasize integrating specialized vocabularies into pre-trained LLMs, prioritizing tokenization quality and cost-effectiveness Ahia et al. (2023); Zhang et al. (2023, 2024); Tejaswi et al. (2024). Liu et al. (2023a) propose a model-agnostic approach for adapting extended vocabularies to LLMs by integrating task-specific vocabularies, prioritizing new tokens, and initializing their embeddings using averaged subword representations. Cui et al. (2023) extend Llama’s existing vocabulary by incorporating an additional 20k Chinese tokens, enhancing its ability to understand and generate Chinese text. Chiappe and Lennon develop an adaptive tokenization algorithm that implements a dynamic tokenization dictionary within the Llama model, updating tokens in real-time based on frequency and contextual relevance.
@@
Cross-lingual transfer enables models trained in one language to adapt to others without retraining from scratch. Key adaptation techniques include embedding initialization, transliteration, and vocabulary extension. Jaavid et al. (2024) used transliteration to convert non-Latin languages into Latin scripts for better knowledge transfer. Zhao et al. (2024) trained a model on millions of target-language tokens without vocabulary extension, achieving performance comparable to models trained on billions of tokens. However, tokenization mismatches reduced inference efficiency. Studies by Csaki et al. (2023); Cui et al. (2023); Raffel et al. (2020); Lin et al. (2024) found that vocabulary extension improves performance while reducing computational inefficiencies. Tejaswi et al. (2024) further explored language-specific LLMs, highlighting trade-offs in adaptation for low-resource languages. Their findings emphasize that while vocabulary expansion enhances efficiency, selecting the right base model and vocabulary size is crucial.
@@
Evaluating LLMs requires benchmarking datasets that assess a wide range of capabilities and tasks. For Bangla, most existing datasets focus on standard NLP tasks. The BanglaNLG benchmark dataset Bhattacharjee et al. (2023) addresses this by integrating six distinct datasets designed to evaluate various aspects of natural language generation (NLG) in Bangla, including Machine Translation, Text Summarization, Question Answering, Multi-turn Dialogue, News Headline Generation, and Cross-lingual Summarization. Beyond NLG, the Region-Specific Native-QA dataset Hasan et al. (2024) was developed to assess the question-answering capabilities of leading LLMs, such as GPT-4o, GPT-4, Gemini, Llama-3, and Mistral. By focusing on regionally relevant queries, this dataset ensures that models are tested in real-world Bangla language contexts. For a broader evaluation of LLMs across multiple tasks, BenLLM Kabir et al. (2024) provides the most comprehensive comparison of model performance. This study benchmarks LLMs against other pretrained models using datasets from diverse sources, offering insights into their strengths and limitations across various NLP tasks.
There is a significant lack of benchmarking datasets for evaluating LLMs. To address this gap, our study developed five benchmarking datasets, each designed to assess different capabilities, including world knowledge and commonsense reasoning.

### v3

@@ Abstract
In this paper, we present TituLLMs, the first large pretrained Bangla LLMs, available in 1b and 3b parameter sizes. Due to computational constraints during both training and inference, we focused on smaller models. To train TituLLMs, we collected a pretraining dataset of approximately $\sim 37$ billion tokens. We extended the Llama-3.2 tokenizer to incorporate language- and culture-specific knowledge, which also enables faster training and inference. There was a lack of benchmarking datasets to benchmark LLMs for Bangla. To address this gap, we developed five benchmarking datasets. We benchmarked various LLMs, including TituLLMs, and demonstrated that TituLLMs outperforms its initial multilingual versions. However, this is not always the case, highlighting the complexities of language adaptation. Our work lays the groundwork for adapting existing multilingual open models to other low-resource languages. To facilitate broader adoption and further research, we have made the TituLLMs models and benchmarking datasets publicly available. [^1] https://github.com/hishab-nlp/titulm
@@
Shahriar Kabir Nahin1, Rabindra Nath Nandi1, Sagor Sarker1, Quazi Sarwar Muhtaseem1, Md Kowsher2, Apu Chandraw Shill1, Md Ibrahim1, Mehadi Hasan Menon1, Tareq Al Muntasir1, Firoj Alam3, 1Hishab Singapore Pte. Ltd, Singapore, 2University of Central Florida, USA 3Qatar Computing Research Institute, Qatar {shahriar.nahin, rabindra.nandi, sagor.sarker}@hishab.co, fialam@hbku.edu.qa
@@ 1 Introduction
The rapid advancements in large language models (LLMs) have reshaped the field of AI, showcasing remarkable versatility across numerous tasks Brown et al. (2020); Ouyang et al. (2022); Achiam et al. (2023); Chowdhery et al. (2023a). These models demonstrate not only an ability to perform various NLP tasks but also an intriguing potential for self-assessment and continuous improvement Liu et al. (2023b); Fu et al. (2023); Chiang et al. (2023).
@@
  • We proposed a novel translation techniques (expressive semantic translation) that helps to develop high quality benchmarking dataset.
@@ 2 Pretraining Data
Pretraining data for Bangla is very limited compared to very high quality data available for English and other high resource languages Penedo et al. (2023); Soldaini et al. (2024). Hence, we needed collect pretraining dataset for training TituLLMs. We have compiled a substantial Bangla raw dataset from a diverse range of sources, encompassing both formal and informal linguistic styles. The dataset is primarily derived from three key sources: web documents, books, and synthetically generated text. The synthetically generated data includes translated data, transliterated data, audio transcripted data, etc. An outline of our data collection and preprocessing pipeline is shown in Figure 7. The final high-quality dataset amounts to approximately 268 GB, with 22 GB designated as a validation set, sampled proportionally to maintain the heterogeneity of the data. In total, the corpus contains $\sim 37$ billion tokens, optimized for training and evaluation across various Bangla language processing applications. In Table 4 (in Appendix) and in Figure 3, we report the distribution of tokens for different sources.
@@ 2.3 Synthetic Data
For training the en-bn machine translation (MT) model, we collected open-source translation data from various platforms, including BanglaNMT Hasan et al. (2020a) and Samanantar Ramesh et al. (2022). Furthermore, we generated synthetic bn-en translation pairs using Bangla news sources and Wikipedia as source data, employing Llama-3.1-8B-Instruct  Touvron et al. (2023) for target data generation. We selected Llama for this task due to its superior English-language capabilities. Using this approach, we created a dataset comprising approximately 60 million translation pairs, which we then used to train the NLLB model. We have named this model as Titu-Translator. On our in-house test dataset, the BLEU score is 37.6.
Evaluation using Scalar Quality Metrics (SQM). We have also evaluated our model using SQM Lankford et al. (2022). It is a structured and interpretable human evaluation of machine translation. Table 1 presents the results of human evaluation conducted by two evaluators across five distinct domains: Business, Entertainment, Politics, Tech, and Sport. Among the evaluated models, Titu-Translator demonstrates strong performance, achieving an overall average score of 4.99, which places it among the top-performing models. It performs particularly well in the Tech and Entertainment domains, with scores up to 5.30 from Human-2 and 4.95 from Human-1. While Indic-Trans-2 slightly outperformed Titu-Translator, we chose Titu-Translator for synthetic pretraining due to its lightweight design and alignment with Bangladeshi linguistic styles. Appendix C describes more about the evaluation process.
Once the model has been trained, we have used it to translate a corpus of English news articles [^5] https://www.kaggle.com/datasets/davidmckinley/all-the-news-dataset into Bangla.
Additionally, we added translated data from the Sangraha dataset Khan et al. (2024) to our corpus. The synthetic data coming from the Sangraha dataset is generated with Indic-Trans-2, which is reported to be the top-performing model for English to Bangla translation in their report Gala et al. (2023). It is also reflected in our SQM evaluation.
Model Business Entertainment Politics Tech Sport Average
H1 H2 H1 H2 H1 H2 H1 H2 H1 H2
Titu Translator 4.8 5.3 4.85 5.4 4.8 5.05 4.95 5.3 4.75 4.65 4.99
nllb-600M 4.45 5.1 4.65 5.0 4.75 4.85 4.85 5.05 4.55 4.2 4.75
csebuetnlp-t5 4.55 4.95 4.65 4.8 4.85 4.9 4.6 4.9 4.65 4.65 4.75
Indic-Trans-2 4.95 5.35 4.85 5.35 4.85 5.35 5.0 4.9 5.0 5.05 5.07
Google Translator 4.6 4.75 4.89 4.65 4.65 4.85 4.95 4.8 5.0 4.55 4.77
Table 1: Human (H) evaluation by two annotators across five domains for each translation model.
@@
The use of romanized text is very common in everyday communication for Bangla Fahim et al. (2024). To address this, we have developed a Bangla-to-Romanized dataset by training an NLLB-based (600M-Distilled) model. For model training, we collected transliteration pairs from the Sangraha dataset and generated additional synthetic transliteration pairs using the GPT-4 model Achiam et al. (2023). We then used this dataset to train the NLLB-based transliteration model. The BLEU score for this model is 65.1, as evaluated on an in-house test dataset. We then used this model to create the transliteration dataset by selecting a small subset of collected Bangla Wikipedia articles.
@@
To enhance the model with conversational capabilities, we enriched our dataset by incorporating conversational data. We have crawled topics (e.g., “Rabindranath Tagore’s contributions to Bengali art”) from Wikipedia and Banglapedia on which we generated conversations between two agents. To achieve this we have developed an agentic system where two agents interact with each other on a given topic. In Appendix Table 7, we have provided examples of topic, roles, and a detail of the prompt. An example conversation is provided in Appendix Figure 7. The average number of turns per conversation is 8. In total, we added $\sim$1 million conversations to the dataset.
@@ 2.4 Sangraha Dataset
Additionally, we enriched our dataset by integrating the open-source Sangraha dataset. It is the largest high-quality, cleaned Indic language corpus. We incorporated the Bangla web data portion of the dataset into our training set.
@@ 3.1 Tokenizer Training
We developed a custom tokenizer for Bangla text using Tiktoken [^6] https://github.com/openai/tiktoken, which employs Byte Pair Encoding (BPE) Sennrich et al. (2016) for tokenization. To train this tokenizer, we sampled 48 GB of data from our pretraining Bangla corpus. Additionally, we modified the original Tiktoken codebase to enhance its efficiency and better accommodate the morphological complexities of Bangla. After training multiple tokenizers on the same subset, we merged each newly trained tokenizer with the existing Llama-3.2 tokenizer. This merging process aimed to preserve the strengths of the original tokenizer while integrating domain-specific vocabulary and improving segmentation for Bangla. To evaluate the performance of each merged tokenizer, we computed the average tokens per word (TPW) on a separate 1 GB sample from the original corpus. Table 5, in Appendix, summarizes the TPW values for both the original Llama-3.2 tokenizer and the newly trained tokenizers. A lower TPW generally indicates more efficient segmentation, which reduces sequence lengths and may enhance downstream model performance.
@@ 3.3 Pretraining
After modifying the models, we pre-trained them on our full dataset using LlamaFactory Zheng et al. (2024). Both models were trained with a context length of 4096, with packing enabled for maximum efficiency. Training for one epoch required 1,750 H100 GPU hours.
@@ 4.2 Banchmarking Datasets
We benchmarked TituLLMs alongside other popular LLMs using five newly prepared evaluation datasets. The dataset is composed of multiple subsets of the benchmarking set, including Bangla MMLU (87,869 entries), Piqa BN (17,177 entries), CommonsenseQA BN (10,962 entries), OpenBookQA BN (5,944 entries), and BoolQ BN (1,976 entries). This distribution highlights the significant dominance of the Bangla MMLU subset within the overall evaluation dataset.
Below, we describe the development process for each dataset. Table 2 presents the distribution and splits of each dataset.
Figure 4: Distribution of an benchmarking dataset totaling $\sim 124$ entries.
Bangla MMLU: We curated multiple-choice questions from various open-source educational websites and textbooks, inspired by the original MMLU dataset Hendrycks et al. (2021). The dataset includes multiple-choice questions from different Bangladeshi exams, such as job exams, the Bangladesh Civil Service Exam, and undergraduate admission exams.
Figure 5 provides a detailed breakdown of the Bangla MMLU dataset, which contains 87,869 questions spanning multiple educational categories. These include University Admission (47,394), Higher Secondary (25,437), Job Exams (9,122), Medical Admission (3,764), and Engineering Admission (2,152). The dataset reflects a diverse range of question types relevant to various levels of academic and professional assessments, making it a comprehensive benchmark for evaluating LLMs in Bangla educational contexts.
Figure 5: The Bangla MMLU dataset comprises a total of 87,869 questions distributed across various educational categories: University Admission (47,394), Higher Secondary (25,437), Job Exams (9,122), Medical Admission (3,764), and Engineering Admission (2,152).
CommonsenseQA Bangla (CSQA): We translated the CommonsenseQA dataset Talmor et al. (2019) into Bangla using our custom translation-based approach, Expressive Semantic Translation (EST). This method generates multiple translations for a sentence and iteratively refines them to select the most suitable version. More details on this approach are discussed in Appendix Section D.
@@
Bangla MMLU Manual - 72,944 14,750 175
@@ 5 Results and Discussion
BoolQ: The BoolQ dataset measures the performance of the model for yes/no question-answering in Bangla. TituLLM-1b achieves 0.53 in the 0-shot setting but drops slightly to 0.51 in the 5-shot setting. In contrast, TituLLM-3B moves from 0.53 (0-shot) to 0.54 (5-shot). However, Llama-3b and Qwen-2.5-1.5b have done much better in this task. As the context length for all BoolQ data matched that of a News document, TituLLM’s performance may drop on long contexts. This suggests further pretraining should target longer contexts Chowdhery et al. (2023b); Kaplan et al. (2020).
@@
Performance of Tokenizer: The superior performance of our models in reasoning tasks is mainly an impact of our extended tokenizer. To justify this, we can observe the results of the BongLLaMa models. These models are continual pretrained models with existing open-source Bangla text corpus. If only the dataset could improve the performance then that would be reflected in BongLLaMA models. However, we observe that they are performing similarly to Llama models. To have an interpretation of our extended tokenizer’s performance we can look into Figure 6. The figure shows Llama tokens and TituLLM tokens for a simple sentence in Bangla with two of the most common words. We observe that Llama tokenizer splits the text into character and byte levels. On the other hand, TituLLM tokenizes the sentence into word or subword levels. As a result, TituLLM can deliver more meaningful tokens than Llama for Bangla text. This is an important advantage of TituLLM that not only enables TituLLM to perform better with smaller datasets but also ensures low latency during inference.
@@ 6 Related Work
The evolution of token adaptation in NLP has progressed from linguistic cues and statistical methods Creutz and Lagus (2006); Luong et al. (2013); Zhang et al. (2023) to phrase-level segmentation Koehn et al. (2007, 2003). The rise of deep learning shifted the focus to subword-level segmentation, enhancing the handling of rare words Sennrich et al. (2016); Kudo (2018); Kudo and Richardson (2018). More recent efforts emphasize integrating specialized vocabularies into pre-trained LLMs, prioritizing tokenization quality and cost-effectiveness Ahia et al. (2023); Zhang et al. (2023, 2024); Tejaswi et al. (2024). Liu et al. (2023a) propose a model-agnostic approach for adapting extended vocabularies to LLMs by integrating task-specific vocabularies, prioritizing new tokens, and initializing their embeddings using averaged subword representations. Cui et al. (2023) extend Llama’s existing vocabulary by incorporating an additional 20k Chinese tokens, enhancing its ability to understand and generate Chinese text. Chiappe and Lennon (2024) develop an adaptive tokenization algorithm that implements a dynamic tokenization dictionary within the Llama model, updating tokens in real-time based on frequency and contextual relevance.
@@
Cross-lingual transfer enables models trained in one language to adapt to others without retraining from scratch. Key adaptation techniques include embedding initialization, transliteration, and vocabulary extension. Jaavid et al. (2024) used transliteration to convert non-Latin languages into Latin scripts for better knowledge transfer. Lai et al. (2024) trained a model on millions of target-language tokens without vocabulary extension, achieving performance comparable to models trained on billions of tokens. However, tokenization mismatches reduced inference efficiency. Studies by Csaki et al. (2023); Cui et al. (2023); Raffel et al. (2020); Lin et al. (2024) found that vocabulary extension improves performance while reducing computational inefficiencies. Tejaswi et al. (2024) further explored language-specific LLMs, highlighting trade-offs in adaptation for low-resource languages. Their findings emphasize that while vocabulary expansion enhances efficiency, selecting the right base model and vocabulary size is crucial.
@@
Evaluating LLMs requires benchmarking datasets that assess a wide range of capabilities and tasks. For Bangla, most existing datasets focus on standard NLP tasks. The BanglaNLG benchmark dataset Bhattacharjee et al. (2023) addresses this by integrating six distinct datasets designed to evaluate various aspects of natural language generation (NLG) in Bangla, including Machine Translation, Text Summarization, Question Answering, Multi-turn Dialogue, News Headline Generation, and Cross-lingual Summarization. Beyond NLG, the Region-Specific Native-QA dataset Hasan et al. (2025) was developed to assess the question-answering capabilities of leading LLMs, such as GPT-4o, GPT-4, Gemini, Llama-3, and Mistral. By focusing on regionally relevant queries, this dataset ensures that models are tested in real-world Bangla language contexts. For a broader evaluation of LLMs across multiple tasks, BenLLM Kabir et al. (2024) provides the most comprehensive comparison of model performance. This study benchmarks LLMs against other pretrained models using datasets from diverse sources, offering insights into their strengths and limitations across various NLP tasks. Other relevant benchmarking efforts include sentiment analysis Hasan et al. (2024), question answering Hasan et al. (2025); Shafayat et al. (2024); Alam et al. (2025), summarization Tanjila et al. (2025), and cultural understanding Kabir et al. (2025).
There remains a notable scarcity of benchmarking datasets for evaluating the emergent capabilities of LLMs, such as world knowledge and cognitive reasoning. To address this gap, we introduce five benchmarking datasets, each designed to assess specific competencies, including world knowledge and commonsense reasoning.